OpenStudy (anonymous):

Write an equation of a line in slope intercept form that is parallel to the line x + 4y = 6 and passes through (-8, 5). Please describe step by step! thanks!

OpenStudy (anonymous):

Ok. So first, you need to put that original line into something we can use. Isolate the y variable. x+4y=6 4y=6-x y=6/4 - x/4 We can rewrite this as: y=-.25 x+ 3/2 Now it's in slope intercept form! So the slope is -.25, and since it's parallel, the slope of the line you're trying to find must be the same. We now have enough info to find the slope-intercept form of your line. y-y1 = m(x - x1) y=m(x - x1) + y1 y = -.25 ( x - (-8)) + 5 y = -.25x - 2 + 5 y = -.25x +3 And that's it! You can check it using this online calculator, if you'd like. https://www.desmos.com/calculator

OpenStudy (anonymous):

This should be in Mathematics By the way ..please stay on topic :)